OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work. -4sqrtx+9=20

OpenStudy (anonymous):

is it \[-4\sqrt{x+9}=20\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so, there is no solution because the left hand side \(-4\sqrt{x+9}\) is negative, but \(20\) is positive

OpenStudy (anonymous):

ok thak you